OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7.

OpenStudy (ivettef365):

to be parallel it needs to have the same slope

OpenStudy (ivettef365):

so you know that the slope is 1, formula y = mx + b, where m = slope

OpenStudy (ivettef365):

now find value of b, by substituting the point (6,-2) on the formula

OpenStudy (anonymous):

I don't know how....

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

you have formula y = mx + b , where y = -2, x =6 and m = 1, find b

OpenStudy (anonymous):

I still don't understand.

OpenStudy (uri):

I think the formula is y = mx + c

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

the letter are irrelevant to find the equation of the line anyways, you have -2 = (1)(6) + b can you solve for b ??

OpenStudy (anonymous):

I litterally don't understand how to do anything... @ivettef365

OpenStudy (ivettef365):

that is algebra, -2 = 6 + b -6 -6 ------------- -8 = b

OpenStudy (ivettef365):

so since you know the value of be you can write the equation of the line as y = x -8

OpenStudy (anonymous):

So is that the answer ? @ivettef365

OpenStudy (ivettef365):

yes, the question ask for the equation of the line, that passes through those points

OpenStudy (anonymous):

Which of the following is the slope between the points (0, -10) and (-4, 0)? Do you understand this one ? @ivettef365

OpenStudy (ivettef365):

you use the slope formula y2 - y1 m = ---------- x2 - x1

OpenStudy (ivettef365):

you have x1, y1 x2, y2 ( 0, -10) (-4, 0)

OpenStudy (anonymous):

Now what ? @ivettef365 :ooooo This is hard -.-

OpenStudy (ivettef365):

just plug in on formula

OpenStudy (ivettef365):

0 - (-10) ------- -4 - 0

OpenStudy (ivettef365):

what do you get?

OpenStudy (anonymous):

NOTHING in this chapter... lol @ivettef365

OpenStudy (ivettef365):

I mean what do you get on that equation that I wrote

OpenStudy (anonymous):

Nothing @ivettef365

OpenStudy (ivettef365):

that is equal to 10/-4 after you simplify it is equal to -5/2 that is the slope

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6 ?

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

to be perpendicular the slope needs to be opposite reciprocal you have slope of -1, opposite reciprocal would be 1

OpenStudy (ivettef365):

then you have y = mx + b, where y = -2 and x = -4, now you try to solve for b

OpenStudy (anonymous):

I dot know how @ivettef365

OpenStudy (ivettef365):

try, you have y = mx + b -2=(1)(-4) + b

OpenStudy (ivettef365):

do you see what I did ???

OpenStudy (anonymous):

noo @ivettef365

OpenStudy (ivettef365):

the formula is y = mx + b I substitute the y for -2, the m for 1 and the x for -4, can you see that

OpenStudy (anonymous):

yes @ivettef365

OpenStudy (ivettef365):

ok now solve for b, you have -2 = 4 + b -4 -4 ------------- -6 = b

OpenStudy (ivettef365):

now that you now the slope = 1 and b = -6 can you write the equation ?

OpenStudy (anonymous):

noo @ivettef365

OpenStudy (ivettef365):

you have y = mx + b , wher (x,y) could be any points, that is why you don't give a value to it, therefore the equation of the line would be y = x -6 (where 1 is the slope, x = 1x and -6 is the b which is the y-int

OpenStudy (anonymous):

Sothat the answer ? @ivettef365

OpenStudy (ivettef365):

yes

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y = 1/2x + 3

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

what is the slope

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

very good, so now what is the opposite reciprocal of 1/2

OpenStudy (anonymous):

2/1 ? @ivettef365

OpenStudy (ivettef365):

Very good, but it is -2/1 which is equal to -2

OpenStudy (ivettef365):

now that you know the slope to be perpendicular let use the formula y = mx + b, you have points (-3,-1) and you know slope is -2 so how would you write it

OpenStudy (anonymous):

-1=-2X+-3 ? @ivettef365

OpenStudy (ivettef365):

almost, y = mx + b y= -1, x=-3, and m = -2 (slope) b is what we need to find so try again

OpenStudy (anonymous):

-1=-2X... IM LOST @ivettef365